OpenStudy (anonymous):

Let f(x) = x + 9 and g(x) = x – 3. What is f(x) • g(x)?

OpenStudy (anonymous):

These are plug-n-chug type questions. f(6) = g(√(6+3)) = g(√9) = g(3) = 7. Using the chain rule f'(x) = g'(√(x+3)) d/dx [√(x+3)] The derivative of √(x + 3) is 1/[2√(x+3)], so f'(x) = 1/[2√(x+3)] g'(√(x+3)) f'(6) = 1/[2√(6+3)] g'(√(6+3)) = 1/(2√9) g' (√9) = (1/6) (8) = 4/3.

OpenStudy (anonymous):

(x+9)*(x-3)=x^2+6x-27